OpenStudy (anonymous):

please help. fan & medal. Use summation notation to write the series 2 + 4+6 +8 + … for 10 terms. In each of these images, the lower limit of the summation notation is either "n=1" or "n = 0"

OpenStudy (anonymous):

what expression can be used to define any even integer?

OpenStudy (anonymous):

\[\sum_{n=1}^{10}2n\]

OpenStudy (anonymous):

2+4+6+..... it is an A.P. a=2,d=4-2=2 tn =a+(n-1) d =2+(n-1)*2 =2+2n-2 =2n \[S10=\sum_{n=1}^{10}\left( 2n \right)\]